• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Markowitz

    Log entry time 00:07:45 on June16,2002

    Entry number 86111

    keyword=Shift summary

    16:00 start of shift, beam gone. MCC talks to our RC. Our currents today have been unacceptably low for the experiment. MCC has called in optics on call, and is going to do a Hall A aperture test. They ask if they can mask ep-ION chambers. Saha does not answer pager, so we have to decline the request.

    12C target is in for this. MCC found nothing and is trying to restore the maximum current achievable to the Hall in the meantime. Our RC defined 65 uA as the minimum acceptable to the experiment.

    19:23 MCC is tuning still but can only deliver 50 uA. We are going to the 15 cm LD2 to take lower current data while they tune the injector. Run 1521 is the first run of this shift on LD2. [Run 1520 had the 12C target.]

    20:05 Ended run 1521 (450K events). Start run 1522

    21:10 Left Q1 Helium Level Fault. Level less than 20%. Called MCC to check if ESR or CHL was having problems. No, but A/B beams are grossly non-coincident and now cannot get any beam from Hall A laser at all.


    Tech on call and TO are downstairs resetting Q1. Luckily while they were still there, the dipole tripped and they were able to reset both magnets in the same access.

    23:10 steering beam into Hall.
    23:30 ready to start a run but DAQ is hung on VME1 (roc 14). Bob Michaels is coming in now for graveyard shift.